PLEASE DO NOT HANDWRITE ANSWER OR DRAW GRAPH WITH HAND! ( Type Answer & Graph Electronically)
Answer #1

Answer:)

Given that,

aS 12 5、多 Ib euiu Aye

Oppurtunity cost of 1 square = 2 berry

so if squash = 4 pound

berry forgone = 8

so, 24-8 = 10 berry

So, this one i am going to specialize in berry

Kew Neighbou

oppurtunity cost of 1 squash = 1 berry

6 berry produced 2 squash

he has lower oppurtunity cost for squash. So comperature advantage in squash.

He will specialize in squash.

Trade:

Let us exchange 1.5 berry for 1 squash

i will produce only berry = 24

the producer in squash = 8

if i consume 16 berry and berry & give away 8 then, 1 will get =1/1.5*8 = 5.33 squash (instead of 4)

New consumption point = A (outside PPF)

Similarly, neighbour produce 8 squash

if he consumer 2 and gives away 6. He will get is 1.5*6= 9.0 berry.

He can consume at point B(outside PPfF

Both gain from trade

Hence Proved.

sorry about the diagram, present i don't option to draw electronically.